### v1

@@ 1 Introduction
One of the standard criteria for correctness in approximate algorithms is the Probably Approximately Correct (PAC) criterion, also referred to as the $(\varepsilon,\delta)$-correct criterion. In our context, a PAC approximate model counter is a randomized algorithm that receives a triple $(F,\varepsilon,\delta)$, where $F$ is an input formula, $\varepsilon>0$ is a tolerance parameter, and $\delta\in(0,1]$ is a confidence parameter, and returns a PAC approximation of the model count of $F$. That is, with the probability at least $1-\delta$ (probably correct), the output should be within a factor of $1\pm\varepsilon$ of the exact model count of $F$ (approximately correct).
@@
Since its initial proposal in 2013 [ChakrabortyMV13], researchers have shown sustained interest in improving the performance of hashing-based techniques [EGSS13b, EGSS13a, CFMS+14, IMMV16, MVCF+15, CMMV16, SM19, AM20, YM21, YCM22]. The most recent update to $\mathsf{ApproxMC}$ [YangM23] significantly improved its efficiency with a novel rounding technique, resulting in the sixth version, $\mathsf{ApproxMC6}$.
@@ 2 Preliminaries
As the name of hashing-based algorithm suggests, ApproxMC uses hash functions as random seeds. For each $n,m\in\mathbb{N}$ with $m\leq n$, we assume access to a hash family $\mathcal{H}(n,m)$ that consists of functions of the form $h:2^{n}\to 2^{m}$. Since its second version [ChakrabortyMV16], ApproxMC exploits a relationship between $\mathcal{H}(n,n)$ and $\mathcal{H}(n,m)$ with $m<n$ that is characterized by prefix-slices. Here, for a given $\vec{x}=(x_{1},\ldots,x_{m},\ldots,x_{n})\in 2^{n}$ and $1\leq m<n$, the $m^{\text{th}}$ prefix-slice $\vec{x}^{(m)}$ of $\vec{x}$ is defined as $\vec{x}^{(m)}=(x_{1},\ldots,x_{m})\in 2^{m}$; and the $m^{\text{th}}$ prefix-slice of a hash function $h:2^{n}\to 2^{n}$ is defined as a function $h^{(m)}:2^{n}\to 2^{m}$ such that $h^{(m)}(y)=h(y)^{(m)}$. Then $\mathcal{H}(n,m)$ is characterized as the set of all $m^{\text{th}}$ prefix-slices of $h\in\mathcal{H}(n,n)$, i.e., $\mathcal{H}(n,m)=\{h^{(m)}\mid h\in\mathcal{H}(n,n)\}$. In this paper, we assume $\mathcal{H}(n,m)$ with $m<n$ is always of this form.
@@
Each hash function $h\in\mathcal{H}(n,m)$ splits the set $2^{n}$ into $2^{m}$ disjoint subspaces, namely $\{h^{-1}(\alpha)\mid\alpha\in 2^{m}\}$ (observe $h^{-1}(\alpha)$ can be empty). Such subspaces are called cells. For a given CNF formula $F$ with $|\mathsf{Vars}(F)|=n$, a hash function $h:2^{n}\to 2^{m}$ and $\alpha\in 2^{m}$, the set of solutions of $F$ that belong to the cell $h^{-1}(\alpha)$ is denoted by $\mathsf{Cell}_{\langle F,h,\alpha\rangle}$; that is, we let $\mathsf{Cell}_{\langle F,h,\alpha\rangle}:=\mathsf{sol}(F)\cap h^{-1}(\alpha)$. We are mostly interested in the cardinality of $\mathsf{Cell}_{\langle F,h,\alpha\rangle}$ for randomly chosen $h$ and $\alpha$, which are generated from $h^{\prime}\in\mathcal{H}(n,n)$ and $\alpha^{\prime}\in 2^{n}$ by taking their $m^{\text{th}}$ prefix-slices; so we mostly omit them from the notation and write $\mathsf{Cell}_{\langle F,m\rangle}$. We also write $\mathsf{Cnt}_{\langle F,m\rangle}$ to denote the value $|\mathsf{Cell}_{\langle F,m\rangle}|$.
@@
Observe, for a given $h\in\mathcal{H}(n,n)$, the sequence of its prefix-slices $h^{(1)},\ldots,h^{(n)}$ can be seen as an iterative splitting policy of the assignment space $2^{n}$: That is, $2^{n}$ is divided into $(h^{(1)})^{-1}(0)$ and $(h^{(1)})^{-1}(1)$; then $(h^{(1)})^{-1}(x)$ is divided into $(h^{(2)})^{-1}(x,0)$ and $(h^{(2)})^{-1}(x,1)$ for each $x\in\{0,1\}$, and so on. Then each $\alpha\in 2^{n}$ naturally induces a sequence of cells $(h^{(1)})^{-1}(\alpha^{(1)})\supseteq\ldots\supseteq(h^{(n)})^{-1}(\alpha^{(n)})$, which can be seen as a sequence of decisions on which cell to take under the iterative splitting policy $h$. In particular, for any fixed $h\in\mathcal{H}(n,n)$, $\alpha\in 2^{n}$ and $1\leq m^{\prime}<m\leq n$, we have $\mathsf{Cell}_{\langle F,m\rangle}\subseteq\mathsf{Cell}_{\langle F,m^{\prime}\rangle}$, and thus $\mathsf{Cnt}_{\langle F,m\rangle}\leq\mathsf{Cnt}_{\langle F,m^{\prime}\rangle}$.
@@
The core counter $\mathsf{ApproxMC6Core}$ works as follows111 The original algorithm in [YangM23] always round up the value of $\mathsf{Cnt}_{\langle F,m\rangle}$ in Line 7 when $\varepsilon\geq 3$. We find this arrangement is not necessary in our argument, so we omit it. . It first samples a hash function $h$ and a string $\alpha$ randomly. After a sanity check (Line 3-4), it finds the smallest number $m$ such that $\mathsf{Cnt}_{\langle F,m\rangle}<\mathsf{thresh}$, meaning that $\mathsf{Cell}_{\langle F,m\rangle}$ is “small” (Line 5). Then it computes $\mathsf{Cnt}_{\langle F,m\rangle}$, possibly “round” it by $\mathsf{rnd}$, and return it after multiplying it by $2^{m}$. Here, $\mathsf{LogSATSearch}$ [ChakrabortyMV16] efficiently finds the number $m$ via a combination of linear and galloping searches. The subroutine $\mathsf{BoundedSAT}$ employs a SAT solver to enumerate distinct solutions of $F$ until either it enumerates all solutions, or the number of enumerated solutions reaches $\mathsf{thresh}$. Notice that $\mathsf{BoundedSAT}$ makes up to $\mathsf{thresh}$ times of SAT calls, and thus, the number of calls to $\mathsf{BoundedSAT}$ is a determining factor of the performance of $\mathsf{ApproxMC6}$. The rounding operation (Line 7) is the new feature of the latest update [YangM23]; it lets us to derive a tighter probability bounds $p_{L}$ and $p_{U}$, and thus offers a smaller repetition number $t$. Apart from that, the high-level structure of $\mathsf{ApproxMC6Core}$ remains the same as the initial version of the core counter.
@@ 4 Technical Details and Soundness
In what follows, we give various statements that involve events in the $\mathsf{ApproxMC6Core}$ (cf. an overview on the bounding argument in §4). Here we formally define them. Recall, for an input $(F,\mathsf{thresh},\mathsf{rnd})$ with $|\mathsf{Vars}(F)|=n$, $\mathsf{ApproxMC6Core}$ randomly samples a hash function $h\in\mathcal{H}(n,n)$ and a vector $\alpha\in 2^{n}$; therefore, we identify an event with the set of all tuples $(h,\alpha)\in\mathcal{H}(n,n)\times 2^{n}$ that trigger the event in $\mathsf{ApproxMC6Core}$. The probability $\mathsf{Pr}[E]$ of an event $E$ is then canonically understood as $\mathsf{Pr}[E]=\frac{|E|}{|\mathcal{H}(n,n)|\times 2^{n}}$.
@@
On events under $F$ with $|\mathsf{Vars}(F)|=n$, we let $L_{i}^{\mathsf{cnt}},U_{i}^{\mathsf{cnt}},L_{i}^{\mathsf{rnd}}$, $U_{i}^{\mathsf{rnd}}$ be all empty for $i\not\in\{1,\ldots,n\}$; we also let $T_{i}$ be empty for $i<1$, and be the full set $\mathcal{H}(n,n)\times 2^{n}$ for $i>n$. This is for technical convenience; it makes the definition of cut-off arguments well-defined even if some cut-off points do not belong to $\{1,\ldots,n\}$ (see (5) and (6)).
@@ 6 Experiments
Recall both $\mathsf{FlexMC}$ and $\mathsf{ApproxMC6}$ are given as Algorithm 1, where the only difference444 We ran experiment with $\varepsilon<3$, and thus the difference in Footnote 1 did not occur. between them is in Line 1: While $\mathsf{FlexMC}$ computes $\mathsf{setParameters}(\varepsilon,\delta)$ by Algorithm 3, $\mathsf{ApproxMC6}$ returns a precomputed value depending on the value of $\varepsilon$ (see [YangM23]). We used a simple ternary search algorithm to realize the FindOptIter procedure in Algorithm 3, whose pseudocode is given in Appendix 0.D. Runtime overhead by Algorithm 3 was less than 1 second.
@@ Appendix 0.D Supplemental Materials on Experiments
For each fixed $\mathsf{thresh}$, we can perform a ternary search on $a_{U}$ (i.e., Algorithm 4) to find the minimum $t$. This is based on our experimental observation that $t$ exhibits unimodal behavior555 We performed an empoirical test with multiple $\varepsilon$, $\delta$, and $\mathsf{thresh}$ (e.g., $\varepsilon=0.8,\delta=0.001,35\leq\mathsf{thresh}\leq 500$ and $\varepsilon=0.4,\delta=0.001,35\leq\mathsf{thresh}\leq 500$). Within this, we didn’t observe any non-unimodal behavior. with respect to $a_{U}$ under a fixed $\mathsf{thresh}$ (see Figure 2). During the ternary search process, for fixed $\mathsf{thresh}$ and $a_{U}$, we employ subroutine Calc_t to determine the minimum $t$. Notably, within this subroutine, a binary search can be applied to find the minimum $t$, enabled by the monotonicity of the error rate with respect to $t$.

### v2

@@ 1 Introduction
One of the standard criteria for correctness in approximate algorithms is the Probably Approximately Correct (PAC) criterion, also referred to as the $(\varepsilon,\delta)$-correct criterion. In our context, a PAC approximate model counter is a randomized algorithm that receives a triple $(F,\varepsilon,\delta)$, where $F$ is an input formula, $\varepsilon>0$ is a tolerance parameter, and $\delta\in(0,1]$ is a confidence parameter, and returns a PAC approximation of the model count of $F$. That is, with the probability at least $1-\delta$ (probably correct), the output should be within a factor of $(1+\varepsilon)^{\pm 1}$ of the exact model count of $F$ (approximately correct).
@@
Since its initial proposal in 2013 [ChakrabortyMV13], researchers have shown sustained interest in improving the performance of hashing-based techniques [EGSS13b, EGSS13a, CFMS+14, IMMV16, MVCF+15, CMMV16, SM19, AM20, YM21, YCM22]. The most recent update111 We take $\mathsf{ApproxMC6}$ [YangM23] as our baseline, as it is the latest version that works for any $\varepsilon$ and $\delta$. A newer version [PoteM025] exists, but it supports only $\varepsilon>1$; see also §7. to $\mathsf{ApproxMC}$ [YangM23] significantly improved its efficiency with a novel rounding technique, resulting in the sixth version, $\mathsf{ApproxMC6}$.
@@ 2 Preliminaries
As the name of hashing-based algorithm suggests, ApproxMC uses hash functions as random seeds. For each $n,m\in\mathbb{N}$ with $m\leq n$, we assume access to a hash family $\mathcal{H}(n,m)$ that consists of functions of the form $h:\{0,1\}^{n}\to\{0,1\}^{m}$. Since its second version [ChakrabortyMV16], ApproxMC exploits a relationship between $\mathcal{H}(n,n)$ and $\mathcal{H}(n,m)$ with $m<n$ that is characterized by prefix-slices. Here, for a given $\vec{x}=(x_{1},\ldots,x_{m},\ldots,x_{n})\in\{0,1\}^{n}$ and $1\leq m<n$, the $m^{\text{th}}$ prefix-slice $\vec{x}^{(m)}$ of $\vec{x}$ is defined as $\vec{x}^{(m)}=(x_{1},\ldots,x_{m})\in\{0,1\}^{m}$; and the $m^{\text{th}}$ prefix-slice of a hash function $h:\{0,1\}^{n}\to\{0,1\}^{n}$ is defined as a function $h^{(m)}:\{0,1\}^{n}\to\{0,1\}^{m}$ such that $h^{(m)}(y)=h(y)^{(m)}$. Then $\mathcal{H}(n,m)$ is characterized as the set of all $m^{\text{th}}$ prefix-slices of $h\in\mathcal{H}(n,n)$, i.e., $\mathcal{H}(n,m)=\{h^{(m)}\mid h\in\mathcal{H}(n,n)\}$. In this paper, we assume $\mathcal{H}(n,m)$ with $m<n$ is always of this form.
@@
Each hash function $h\in\mathcal{H}(n,m)$ splits the set $\{0,1\}^{n}$ into $2^{m}$ disjoint subspaces, namely $\{h^{-1}(\alpha)\mid\alpha\in\{0,1\}^{m}\}$ (observe $h^{-1}(\alpha)$ can be empty). Such subspaces are called cells. For a given CNF formula $F$ with $|\mathsf{Vars}(F)|=n$, a hash function $h:\{0,1\}^{n}\to\{0,1\}^{m}$ and $\alpha\in\{0,1\}^{m}$, the set of solutions of $F$ that belong to the cell $h^{-1}(\alpha)$ is denoted by $\mathsf{Cell}_{\langle F,h,\alpha\rangle}$; that is, we let $\mathsf{Cell}_{\langle F,h,\alpha\rangle}:=\mathsf{sol}(F)\cap h^{-1}(\alpha)$. We are mostly interested in the cardinality of $\mathsf{Cell}_{\langle F,h,\alpha\rangle}$ for randomly chosen $h$ and $\alpha$, which are generated from $h^{\prime}\in\mathcal{H}(n,n)$ and $\alpha^{\prime}\in\{0,1\}^{n}$ by taking their $m^{\text{th}}$ prefix-slices; so we mostly omit them from the notation and write $\mathsf{Cell}_{\langle F,m\rangle}$. We also write $\mathsf{Cnt}_{\langle F,m\rangle}$ to denote the value $|\mathsf{Cell}_{\langle F,m\rangle}|$.
@@
Observe, for a given $h\in\mathcal{H}(n,n)$, the sequence of its prefix-slices $h^{(1)},\ldots,h^{(n)}$ can be seen as an iterative splitting policy of the assignment space $\{0,1\}^{n}$: That is, $\{0,1\}^{n}$ is divided into $(h^{(1)})^{-1}(0)$ and $(h^{(1)})^{-1}(1)$; then $(h^{(1)})^{-1}(x)$ is divided into $(h^{(2)})^{-1}(x,0)$ and $(h^{(2)})^{-1}(x,1)$ for each $x\in\{0,1\}$, and so on. Then each $\alpha\in\{0,1\}^{n}$ naturally induces a sequence of cells $(h^{(1)})^{-1}(\alpha^{(1)})\supseteq\ldots\supseteq(h^{(n)})^{-1}(\alpha^{(n)})$, which can be seen as a sequence of decisions on which cell to take under the iterative splitting policy $h$. In particular, for any fixed $h\in\mathcal{H}(n,n)$, $\alpha\in\{0,1\}^{n}$ and $1\leq m^{\prime}<m\leq n$, we have $\mathsf{Cell}_{\langle F,m\rangle}\subseteq\mathsf{Cell}_{\langle F,m^{\prime}\rangle}$, and thus $\mathsf{Cnt}_{\langle F,m\rangle}\leq\mathsf{Cnt}_{\langle F,m^{\prime}\rangle}$.
@@
The core counter $\mathsf{ApproxMC6Core}$ works as follows222 The original algorithm in [YangM23] always round up the value of $\mathsf{Cnt}_{\langle F,m\rangle}$ in Line 7 when $\varepsilon\geq 3$. We find this arrangement is not necessary in our argument, so we omit it. . It first samples a hash function $h$ and a string $\alpha$ randomly. After a sanity check (Line 3-4), it finds the smallest number $m$ such that $\mathsf{Cnt}_{\langle F,m\rangle}<\mathsf{thresh}$, meaning that $\mathsf{Cell}_{\langle F,m\rangle}$ is “small” (Line 5). Then it computes $\mathsf{Cnt}_{\langle F,m\rangle}$, possibly “round” it by $\mathsf{rnd}$, and return it after multiplying it by $2^{m}$. Here, $\mathsf{LogSATSearch}$ [ChakrabortyMV16] efficiently finds the number $m$ via a combination of linear and galloping searches. The subroutine $\mathsf{BoundedSAT}$ employs a SAT solver to enumerate distinct solutions of $F$ until either it enumerates all solutions, or the number of enumerated solutions reaches $\mathsf{thresh}$. Notice that $\mathsf{BoundedSAT}$ makes up to $\mathsf{thresh}$ times of SAT calls, and thus, the number of calls to $\mathsf{BoundedSAT}$ is a determining factor of the performance of $\mathsf{ApproxMC6}$. The rounding operation (Line 7) is the new feature of the latest update [YangM23]; it lets us to derive a tighter probability bounds $p_{L}$ and $p_{U}$, and thus offers a smaller repetition number $t$. Apart from that, the high-level structure of $\mathsf{ApproxMC6Core}$ remains the same as the initial version of the core counter.
@@ 4 Technical Details and Soundness
In what follows, we give various statements that involve events in the $\mathsf{ApproxMC6Core}$ (cf. an overview on the bounding argument in §4). Here we formally define them. Recall, for an input $(F,\mathsf{thresh},\mathsf{rnd})$ with $|\mathsf{Vars}(F)|=n$, $\mathsf{ApproxMC6Core}$ randomly samples a hash function $h\in\mathcal{H}(n,n)$ and a vector $\alpha\in\{0,1\}^{n}$; therefore, we identify an event with the set of all tuples $(h,\alpha)\in\mathcal{H}(n,n)\times\{0,1\}^{n}$ that trigger the event in $\mathsf{ApproxMC6Core}$. The probability $\mathsf{Pr}[E]$ of an event $E$ is then canonically understood as $\mathsf{Pr}[E]=\frac{|E|}{|\mathcal{H}(n,n)|\times 2^{n}}$.
@@
On events under $F$ with $|\mathsf{Vars}(F)|=n$, we let $L_{i}^{\mathsf{cnt}},U_{i}^{\mathsf{cnt}},L_{i}^{\mathsf{rnd}}$, $U_{i}^{\mathsf{rnd}}$ be all empty for $i\not\in\{1,\ldots,n\}$; we also let $T_{i}$ be empty for $i<1$, and be the full set $\mathcal{H}(n,n)\times\{0,1\}^{n}$ for $i>n$. This is for technical convenience; it makes the definition of cut-off arguments well-defined even if some cut-off points do not belong to $\{1,\ldots,n\}$ (see (5) and (6)).
@@ 6 Experiments
Recall both $\mathsf{FlexMC}$ and $\mathsf{ApproxMC6}$ are given as Algorithm 1, where the only difference555 We ran experiment with $\varepsilon<3$, and thus the difference in Footnote 2 did not occur. between them is in Line 1: While $\mathsf{FlexMC}$ computes $\mathsf{setParameters}(\varepsilon,\delta)$ by Algorithm 3, $\mathsf{ApproxMC6}$ returns a precomputed value depending on the value of $\varepsilon$ (see [YangM23]). We used a simple ternary search algorithm to realize the FindOptIter procedure in Algorithm 3, whose pseudocode is given in Appendix 0.D. Runtime overhead by Algorithm 3 was less than 1 second.
@@ 7 Related Works
Yet another new variant of $\mathsf{ApproxMC}$, called $\mathsf{ApproxMC7}$, has been released rather recently [PoteM025]. The experiments in the paper demonstrate better runtime performance of $\mathsf{ApproxMC7}$ over $\mathsf{ApproxMC6}$; however, $\mathsf{ApproxMC7}$ requires the assumption $\varepsilon>1$ for its correctness. Extending our formalism to include $\mathsf{ApproxMC7}$ as a particular instance is left for future work. Indeed, $\mathsf{ApproxMC7}$ resembles $\mathsf{ApproxMC6}$ except for the different choice of internal parameters—for example, its core counter can be viewed as $\mathsf{ApproxMC6Core}$ with specific values of $\mathsf{thresh}$ and rounding parameters. However, it sets $\mathsf{thresh}=1$, which falls outside the range of our analysis; there are also subtle differences in the derivation of the error probability bounds. Because of these, our current formalism cannot instantiate $\mathsf{ApproxMC7}$.
@@ 8 Conclusion
Acknowledgment. We thank Mate Soos and Jiong Yang for their help in finding the necessary resources for our experiments.
@@ Appendix 0.D Supplemental Materials on Experiments
For each fixed $\mathsf{thresh}$, we can perform a ternary search on $a_{U}$ (i.e., Algorithm 4) to find the minimum $t$. This is based on our experimental observation that $t$ exhibits unimodal behavior666 We performed an empoirical test with multiple $\varepsilon$, $\delta$, and $\mathsf{thresh}$ (e.g., $\varepsilon=0.8,\delta=0.001,35\leq\mathsf{thresh}\leq 500$ and $\varepsilon=0.4,\delta=0.001,35\leq\mathsf{thresh}\leq 500$). Within this, we didn’t observe any non-unimodal behavior. with respect to $a_{U}$ under a fixed $\mathsf{thresh}$ (see Figure 2). During the ternary search process, for fixed $\mathsf{thresh}$ and $a_{U}$, we employ subroutine Calc_t to determine the minimum $t$. Notably, within this subroutine, a binary search can be applied to find the minimum $t$, enabled by the monotonicity of the error rate with respect to $t$.